OpenStudy (anonymous):

If the cost of manufacturing x items is x^3 + 21x^2 + 110x + 20 , find the marginal cost at x = 100 and the actual cost of manufacturing the 100th item.

OpenStudy (anonymous):

I did it but i got wrong C(x) = x^3 + 21x^2 + 110x + 20 F’9x) = 3x^2 + 42 _ 110 100^3 + 21(100)^2 + 110(100) + 20 = 1221020 99^3 + 21(99)^2 + 110(99) + 20 = 1187030 c(100) – c(99) = 1033990 f’(x) = 3x^2 + 42 + 110 = 24310 actual cost : 1033990 marginal cost: 34310

OpenStudy (anonymous):

My teacher told me to re evaluate c(100) - c(99)

OpenStudy (anonymous):

@tcarroll010 can u help?

OpenStudy (anonymous):

c(100) is correct, so is c(99) c(100) - c(99) = 1221020 - 1187030 = 34020 - 30 = 33990

OpenStudy (anonymous):

ok